# Discovery of energy landscapes towards optimized quantum transport: Environmental effects and long-range tunneling

### v1

@@ III.3 OQS Model II
To probe temperature-induced effects, we next study the three-site model at finite temperature, as described by OQS Model II, Eq. (6). Results are presented in Fig. 4, and we make the following observations: (i) When $J_{2}=0$, the optimal energy profile is no longer flat. Instead, there is a preference for a structured energy profile, with $\varepsilon_{2}$ elevated above the other two sites and $\varepsilon_{3}<\varepsilon_{1}$, see panels (a)-(b). Although at infinite temperature raising the intermediate level provides no advantage, at finite temperature this strategy enhances the flux. One way to rationalize this landscape is to point out that since the intermediate level is high in energy, its steady-state population is low (c), leaving more population available at the exit site 3. (ii) When $J_{2}\neq 0$, the strategy identified in Fig. 3 was to suppress destructive interference by lowering $\varepsilon_{2}$. However, this approach is less effective at intermediate temperatures because it causes carrier accumulation at the middle site (site 2) rather than at the exit site. As such, we find that the structure in Fig. 4 (d)-(e) balances interference effects with thermalization, yielding an optimal profile where $\varepsilon_{2}\sim\varepsilon_{3}<\varepsilon_{1}$.
@@ IV Optimal energy landscapes in long chains
Our goal is to develop physical intuition and offer practical guidelines for identifying energy profiles that enhance carrier transport in quasi-one-dimensional chains. We examine two central factors: (i) the interplay between environmental effects and coherent dynamics, and (ii) the role of long-range tunneling relative to short-range tunneling.
@@
In the following, we present results for nine- and ten-site systems. These cases were selected to investigate possible even-odd effects in the chain length. For completeness, Appendix B provides parallel simulation results for intermediate chain lengths of five and six sites.
As a reminder, for an $N$-site chain, we optimize over the N1N-1 energy levels relative to site 1, which is fixed as a reference point to zero energy. Other parameters are kept constant. We typically set the maximum tunneling energy to $J_{\text{max}}=0.2$ and the leak rate to $\gamma_{l}=0.1$. In OQS Model I, the dephasing rate is set to $\Gamma=0.1$, while in OQS Model II we use a temperature of $T=0.2$ and the dimensionless system-bath coupling $\Gamma_{0}=0.1$.
@@
Figure 6: Converged optimal energy profiles found when optimizing transport without environmental interactions, and their corresponding population flux. Parameters are the same as in Fig. 5.
@@ IV.1 Coherent Model
Fig. 5(a)-(b) show the energy profiles that optimize transport along the chain in the coherent limit. For short-range tunneling ($\alpha=3$) the profiles remain nearly flat, reflecting a uniform landscape. In contrast, the long-range tunneling case ($\alpha=1$) dramatically reshapes the energy landscape by “pushing” intermediate energy levels away from resonance with both the entrance and exit sites, creating a more complex pattern.
This contrast is echoed in the population distributions shown in Fig. 5(c)-(d). Under short-range tunneling, populations are evenly spread across the chain, pointing to ballistic transport. Meanwhile, under long-range tunneling, intermediate sites generally hold low populations, acting as transient bridges, though in the 10-site chain some intermediate levels surprisingly maintain high populations. Generally though, the chain population exemplifies the competition between the classically-required population gradient for transport (see the explanation of ENAQT in [50, 51]), and our problem-specific requirement that the last site population be large so that the population flux is large.
@@
To uncover the underlying transport mechanisms at play, we analyze in more detail the steady-state populations and coherences under both coupling models. The steady-state populations for the short-range tunneling case ($\alpha=3$), shown in pink in Fig. 5(c)-(d), offers insight into why the corresponding energy profile is optimal. The population distribution exhibits a shallow gradient across the chain, indicative of near-ballistic transport, where the carriers flow with minimal scattering or localization. In contrast, the steady-state populations for the long-range tunneling case ($\alpha=1$), shown in green in Fig. 5(c)-(d), reveal a strikingly different transport strategy. Here, population accumulates predominantly at the first and last sites, with suppressed occupation of the intermediate sites. This suggests that the optimal energy profile minimizes the occupation of intermediate sites by energetically detuning them. This configuration helps avoid destructive interference effects, which are most pronounced when levels are near resonance.
@@
The density matrix maps in Fig. 5(e)-(h) corroborate the proposed transport mechanisms, with short-range tunneling ($\alpha=3$) facilitating length-independent ballistic motion. In contrast, a direct tunneling between the entry and exit sites, or possibly assisted by some intermediate levels, is shown in the long range ($\alpha=1$) case. The energetically displaced levels exhibit strong coherence with each other, suggesting that they act as a “bridge” that helps to preserve the carrier’s coherence throughout the process. However, we note that the steady-state populations in the coherent long-range model are highly sensitive to small variations in site energies. We suspect that the energy and population profiles depicted in Fig. 5 correspond to very sharp maxima in the population flux landscape, reflecting narrow “sweet spots” where transport flux peaks.
@@
A natural question that arises is whether the optimized structures we obtain are constrained by our choice of initial optimizer conditions. In other words, are we truly finding the global maximum within the parameter space, or could there be even better solutions hidden beyond our current sampling? In Figs. 6(a)-(d) we present the successfully-optimized profiles; see Appendix A for details on the optimization procedure. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 5(a)-(b). Although the optimization was repeated with many randomized initial “guesses” for optimal profiles, from Fig. 6(a) and (c) we find that in the short-range case, many initial conditions reach the same optimal solution or closely similar solutions and, correspondingly, a similar flux. In contrast, for the long-range model, Fig. 6(b) and (d) show that the parameter space apparently supports many local maxima. We identify several highly efficient solutions alongside a variety of sub-optimal ones. Notably, the best-performing solutions consistently share a key feature: the energies of the intermediate sites are strongly detuned from those of the entry and exit sites.
@@
Fig. 13 in Appendix B shows that five- and six-site chains exhibit similar patterns in energy profiles, population distribution, and coherence behaviors for both short- and long-range tunneling regimes. For short-range tunneling, the energy profile remains nearly uniform across the chain, hovering close to zero under our site 1 energy convention. In contrast, the long-range tunneling scenario favors setting the entrance and exit sites near resonance, while energetically detuning the intermediate sites. Interestingly, the most efficient solutions typically feature intermediate sites shifted to negative energies. However, once again solutions under the long-range model are sensitive; small modifications in parameters can drastically alter the optimal energy landscape. This sensitivity gives rise to many locally-optimal energy profiles with intermediate sites detuned to positive energies, shown in Fig. 14.
@@
We summarize our findings on optimizing carrier transport in chains isolated from environmental effects. We refer to this as the “Coherent model” (CM) design rules: For short-range tunneling, the optimal strategy favors an almost uniform energy landscape that supports near-ballistic transport. In contrast, when long-range tunneling plays a significant role, the optimal energy profile leverages these extended couplings while simultaneously suppressing destructive interference. This results in a non-uniform, corrugated energy landscape that energetically detunes intermediate sites to facilitate more direct transfer between the chain’s endpoints.
@@ IV.2 OQS Model I
We continue our efforts to identify and classify chain configurations that enhance transport under varying conditions. Optimized structures under local dephasing for the $N=9$ and $N=10$ chains are presented in Fig. 7 (a)-(b). Consistent with previous findings, we confirm that no significant even–odd effects are present: the optimal energy profiles are remarkably similar in the nine-site and ten-site chains. In all cases, the first and last sites exhibit comparable energies, consistent with the behavior observed in the short three-site model.
@@
Figure 8: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport in OQS Model I, with the best-performing profile shown in Fig. 7.
Figure 9: Population flux $\eta$ as a function of environmental noise $\Gamma$ for (a) $N=9$-site chain using the optimized energy profile for OQS Model I from Fig. 7(a); (b) $N=10$-site chain using the optimized energy profile for OQS I from Fig. 7(b). In the long range case, the corrugated energy landscape gives rise to an ENAQT effect.
@@
Figure 11: Converged optimal energy profiles and corresponding fluxes obtained from transport optimization under OQS Model II. Parameters are as in Fig. 10(a)-(b).
@@
Previous studies have shown that ENAQT in systems with nearest-neighbor tunneling arises from a competition between dephasing-driven population equalization and the requirement of a population gradient to drive flow from the excitation site to the sink, as classically described by Fick’s law [50, 51]. A pronounced population gradient signals predominantly classical transport behavior. The optimal structures under short-range tunneling, shown in Fig. 7(c)-(d) (orange squares) appear to conform to this classical picture.
To better understand the seemingly nonintuitive energy profiles observed in the long-range tunneling case, Fig. 7(a)-(b), we turn to the corresponding steady-state populations in Fig. 7(c)-(d). The population initially decreases in a roughly linear fashion, then saturates toward the end of the chain (purple triangles). This suggests that quantum effects are shaping the transport behavior under long-range interactions. Remarkably, enabled by the nontrivial energy structure, the system appears to establish a ballistic-like transport regime, evidenced by the nearly constant population towards the end of the chain.
Steady-state coherences in the site basis further shed light on the underlying transport mechanisms. Fig. 7(e)-(h) displays the steady-state coherences for the optimal energy profiles shown in Fig. 7(a)-(b). In the short-range case, significant coherences are observed only between nearest and next-nearest neighbors (Figs. 7(e) and (g)) indicating that carrier transmission occurs predominantly via adjacent sites. In contrast, the long-range tunneling cases (Figs. 7(f) and (h)) exhibit pronounced coherences between entry site 1 and several other remote sites, particularly the exit site $N$. It is interesting to note that site 2 shows only weak coherences with the rest of the chain, which aligns with its significantly lower energy relative to the other sites in the long-range energy profiles [Fig. 7(a)-(b)]. Taken together with the steady-state populations, these findings suggest that, much like in the coherent case, long-range tunneling favors an optimal energy profile that promotes direct transport from the initial to the terminal site. By effectively bypassing the intermediate sites, this configuration enhances long-range transfer and boosts the overall flux.
@@
To complement our analysis and illustrate the related ENAQT behavior under our choice of parameters, we present in Fig. 9 plots of the population flux, $\eta$, as a function of environmental noise. Here, we adopt the structures identified as optimal for short- and long-range chains of nine and ten sites, and study the flux while varying the dephasing rate, $\Gamma$. Notably, a turnover in flux, a hallmark of ENAQT, appears only in systems with long-range tunneling, while the flux in purely short-range-tunneling systems shows no such peak. The peak of the flux in the long-range case is not centered precisely at the $\Gamma$ value used to obtain the optimal energy profiles; optimization was constrained to a specific value of $\Gamma$ (indicated by the dashed line). However, for both nine- and ten-site chains the long-range peaks are close to our chosen $\Gamma$, reflecting that we are working in an ENAQT regime.
In Appendix B, we repeat the optimization process of OQS Model I with shorter chains of five and six sites; see Fig. 15. Our findings for these shorter chains are consistent with the nine- and ten-site chains. For models with short-range tunneling, we see nearly flat energy profiles and a constant gradient population. In contrast, long-range tunneling produces non-monotonic, corrugated energy landscapes, with population profiles exhibiting a clear transition from a decaying trend to an almost constant distribution along the chain.
We summarize our key findings on optimizing carrier transport in chains subject to local dephasing with the following OQS I design rule: When local dephasing is strong and comparable in scale to the tunneling energies, the short-range tunneling model favors classical-like conduction with population decaying with length. This is achieved through nearly flat energy landscapes with only mild detuning from the entry and exit sites. In contrast, when long-range tunneling is allowed, optimal transport arises when the system successfully establishes ballistic-like conduction, even if only partial, as reflected by the steady-state population profiles. This transport regime is enabled by a highly nontrivial energy landscape that suppresses interference and promotes efficient end-to-end transfer.
The observation of a relatively flat population profile at some sites away from the entrance ($n=4$ in Fig. 7(c)-(d) and Fig. 15(c)-(d)) under long-range tunneling is particularly striking. It demonstrates that ballistic-like transport can emerge even in chains with highly non-uniform energy landscapes, provided that the energy structure is carefully balanced with the long-range tunneling amplitudes and environmental effects.

### v2

@@ III.3 OQS Model II
To probe instead temperature-induced effects, we next study the three-site model at finite temperature, as described by OQS Model II, Eq. (6). Results are presented in Fig. 4, and we make the following observations: (i) When $J_{2}=0$, the optimal energy profile is no longer flat. Instead, there is a preference for a structured energy profile, with $\varepsilon_{2}$ elevated above the other two sites and $\varepsilon_{3}<\varepsilon_{1}$, see panels (a)-(b). At infinite temperature, or under OQS Model I, raising the intermediate level provides no advantage. However, at finite temperature under OQS Model II, this strategy enhances the flux. One way to rationalize this landscape is to point out that since the intermediate level is high in energy, its steady-state population is low (c), leaving more population available at the exit site 3. (ii) When $J_{2}\neq 0$, the strategy identified in Fig. 3 was to suppress destructive interference by lowering $\varepsilon_{2}$. However, this approach is less effective at intermediate temperatures because it causes carrier accumulation at the middle site (2), rather than at the exit site. As such, we find that the structure in Fig. 4 (d)-(e) balances interference effects with thermalization, yielding an optimal profile where $\varepsilon_{2}\sim\varepsilon_{3}<\varepsilon_{1}$.
@@ IV Optimal energy landscapes in long chains
Our aim is to build physical intuition and provide practical design principles for energy profiles that enhance carrier transport in quasi-one-dimensional chains. We focus on two key factors: (i) the interplay between coherent dynamics and environmental effects, and (ii) the influence of long-range versus short-range tunneling.
@@
In the following, we present results for nine- and ten-site systems. These cases were selected to investigate possible even-odd effects in the chain length. For completeness, Appendix B provides parallel simulation results for intermediate chain lengths of five and six sites. Given the complexity of the model, we did not attempt to solve it analytically. However, it is worth pointing out that mapping the quantum transport system to classical kinetic networks may allow derivations of closed-form expressions [11].
In addition to considerations of coherent and incoherent effects, the choice of $\gamma_{l}$ further controls this transport trend. However, to fairly compare results between different models, we keep this parameter fixed and instead focus on the roles of the inner-chain energy profile and the environmental model.
As a reminder, for an $N$-site chain, we optimize over the N1N-1 energy levels relative to site 1, which is fixed as a reference point to zero energy. Other parameters are kept constant. We set the maximum tunneling energy to $J_{\text{max}}=0.2$ and the leak rate to $\gamma_{l}=0.1$. In OQS Model I, the dephasing rate is set to $\Gamma=0.1$, while in OQS Model II we use a temperature of $T=0.2$ and set the dimensionless system-bath coupling to $\Gamma_{0}=0.1$.
@@
Figure 6: Converged optimal energy profiles found when optimizing transport without environmental interactions, and their corresponding population flux. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 5(a)-(b). Parameters are the same as in Fig. 5.
@@ IV.1 Coherent Model
Fig. 5(a)-(b) show the energy profiles that optimize transport along the chain in the coherent limit. For short-range tunneling ($\alpha=3$) the profiles remain nearly flat, reflecting an essentially uniform landscape. In contrast, the long-range tunneling case ($\alpha=1$) dramatically reshapes the energy landscape by “pushing” intermediate energy levels away from resonance with both the entrance and exit sites, creating a more complex pattern.
This reflects what we saw for the 3-site system in Eq. (15), where a large, negative middle-site detuning led to greater transport under long-range tunneling ($J_{2}\neq 0$). We intuit that a similar principle governs the transport behavior in longer chains, although in this case we see that the intermediate sites may be positively or negatively detuned.
@@
The contrast between short-range and long-range tunneling is echoed in the population distributions shown in Fig. 5(c)-(d). To uncover the underlying transport mechanisms at play, we analyze in more detail the steady-state populations and coherences under both coupling models. The steady-state populations for the short-range tunneling case ($\alpha=3$), shown in pink in Fig. 5(c)-(d), offer insight into why the corresponding energy profile is optimal. The population distribution exhibits a shallow gradient across the chain, indicative of quasi-ballistic transport, where the carriers flow with minimal scattering or localization. In contrast, the steady-state populations for the long-range tunneling case ($\alpha=1$), shown in green in Fig. 5(c)-(d), reveal a strikingly different transport strategy. Here, population accumulates predominantly at the first and last sites, with suppressed occupation of the intermediate sites. This suggests that the optimal energy profile minimizes the occupation of intermediate sites by energetically detuning them, and helps avoid destructive interference effects, which are most pronounced when levels are near resonance. This proposed explanation on the mitigation of destructive interference effects could be examined in future work by decomposing the total flux into its pathways’ components and analyzing the site-to-site contributions, following the approach of Ref. [67]. Though in the 10-site chain some intermediate levels surprisingly maintain high populations, generally the chain population exemplifies the competition between the classically-required population gradient for transport (see the explanation of ENAQT in [115, 116]), and our problem-specific requirement that the last site population be large so that the population flux is large.
@@
The density matrix maps in Fig. 5(e)-(h) corroborate the proposed transport mechanisms, with short-range tunneling ($\alpha=3$) facilitating transport through site-to-site coherent transport. In contrast, a direct tunneling between the entry and exit sites, or possibly assisted by some intermediate levels, is shown in the long range ($\alpha=1$) case. The energetically-displaced levels exhibit strong coherence with each other, suggesting that they act as a “bridge” that helps to preserve the carrier’s coherence throughout the process. However, we note that the steady-state populations in the coherent long-range model are highly sensitive to small variations in site energies. We suspect that the energy and population profiles depicted in Fig. 5 correspond to very sharp maxima in the population flux landscape, reflecting narrow “sweet spots” where transport flux peaks; for more details, see Appendix A.3.
@@
A natural question that arises is whether the optimized structures we obtain are constrained by our choice of initial optimizer conditions. In other words, are we truly finding the global maximum within the parameter space, or could there be even better solutions hidden beyond our current sampling? In Figs. 6(a)-(d) we present the successfully-optimized profiles; see Appendix A for details on the optimization procedure. Although the optimization was repeated with many randomized initial “guesses” for optimal profiles, from Fig. 6(a) and (c) we find that in the short-range case, many initial conditions reach the same optimal solution or closely similar solutions and, correspondingly, a similar flux. In contrast, for the long-range model, Fig. 6(b) and (d) show that the parameter space apparently supports many local maxima. We identify several highly efficient solutions alongside a variety of sub-optimal ones. Notably, the best-performing solutions consistently share a key feature: the energies of the intermediate sites are strongly detuned from those of the entry and exit sites.
@@
Figure 17 in Appendix B demonstrates that five- and six-site chains display similar patterns in energy profiles, population distributions, and coherence dynamics across both short- and long-range tunneling regimes. For short-range tunneling, the energy profile remains nearly uniform across the chain, hovering close to zero under our site 1 energy convention. In contrast, the long-range tunneling scenario favors setting the entrance and exit sites near resonance, while energetically detuning the intermediate sites. Interestingly, the most efficient solutions typically feature intermediate sites shifted to negative energies. However, once again solutions under the long-range model are sensitive; small modifications in parameters can drastically alter the optimal energy landscape. This sensitivity gives rise to many locally-optimal energy profiles with intermediate sites detuned to positive energies, shown in Fig. 18.
@@
We summarize our findings on optimizing carrier transport in chains isolated from environmental effects. We refer to these as the “Coherent model” (CM) design rules: for short-range tunneling, the optimal strategy favors an almost uniform energy landscape that supports quasi-ballistic transport. In contrast, when long-range tunneling plays a significant role, we propose that the optimal energy profile leverages these extended couplings while simultaneously suppressing destructive interference. This results in a non-uniform, corrugated energy landscape that energetically detunes intermediate sites to facilitate more direct transfer between the chain’s endpoints.
@@ IV.2 OQS Model I
We extend our analysis to identify and classify chain configurations that maximize transport under varying conditions. Optimized structures for $N=9$ and $N=10$ chains under local dephasing are shown in Fig. 7 (a)–(b). Consistent with earlier observations, we find no significant even–odd effects: the optimal energy profiles for nine- and ten-site chains are strikingly similar. In all cases, the first and last sites exhibit comparable energies, mirroring the behavior seen in the minimal three-site model.
@@
To understand why these profiles are better than a uniform energy profile, it is helpful to return to the 3-site system results in Section III. We see from equation (20) that in the three-site system under OQS Model I, a uniform energy profile leads to a decrease in transport with increasing $\Gamma$, when $J_{2}<J_{1}$. A similar effect likely applies to the longer 9- and 10-site chains; hence, it is not surprising that optimal energy profiles are not flat.
Figure 8: Converged optimal energy profiles found when optimizing transport in OQS Model I and their corresponding population flux. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 7(a)-(b). Parameters are the same as in Fig. 7.
Figure 9: Population flux $\eta$ as a function of environmental noise $\Gamma$ for (a) $N=9$-site chain using the optimized energy profile for OQS Model I from Fig. 7(a); (b) $N=10$-site chain using the optimized energy profile for OQS I from Fig. 7(b). In the long range case, the corrugated energy landscape gives rise to an ENAQT effect. The dashed line is the fixed value of $\Gamma$ used for the optimization.
@@
Figure 11: Converged optimal energy profiles found when optimizing transport in OQS Model II and their corresponding population flux. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 10(a)-(b). Parameters are the same as in Fig. 10
@@
Previous studies have shown that ENAQT in systems with nearest-neighbor tunneling arises from a competition between dephasing-driven population equalization and the requirement of a population gradient to drive flow from the excitation site to the sink, as classically described by Fick’s law [115, 116]. A pronounced population gradient signals predominantly classical transport behavior. The optimal structures under short-range tunneling, shown in Fig. 7(c)-(d) (orange squares) appear to conform to this classical picture: the population profile shows that there is nearly a constant population gradient in the chain, driving a diffusive flux, as described by Fick’s law.
To better understand the seemingly non-intuitive energy profiles observed in the long-range tunneling case, Fig. 7(a)-(b), we turn to the corresponding steady-state populations in Fig. 7(c)-(d). The population initially decreases, then saturates toward the end of the chain (purple triangles). Remarkably, the interplay between strong coherent effects enabled by large tunneling elements, and incoherent effects, enabled by comparable dephasing rates leads to a nontrivial energy structure that allows the system to establish an efficient transport, evidenced by the nearly constant population developing halfway through the chain.
Steady-state coherences in the site basis further shed light on the underlying transport mechanisms. Fig. 7(e)-(h) displays the steady-state coherences for the optimal energy profiles shown in Fig. 7(a)-(b). In the short-range case, significant coherences are observed only between nearest and next-nearest neighbors [Figs. 7(e) and (g)] indicating that carrier transmission occurs predominantly via adjacent sites. In contrast, the long-range tunneling cases [Figs. 7(f) and (h)] exhibit pronounced coherences between entry site 1 and several other remote sites, particularly the exit site $N$. It is interesting to note that site 2 shows only weak coherences with the rest of the chain, which aligns with its significantly lower energy relative to the other sites in the long-range energy profiles [Fig. 7(a)-(b)]. Taken together with the steady-state populations, these findings suggest that, much like in the coherent case, long-range tunneling favors an optimal energy profile that promotes direct transport from the initial to the terminal site. By effectively bypassing the intermediate sites, this configuration enhances long-range transfer and boosts the overall flux.
@@
To complement our analysis and illustrate the related ENAQT behavior under our choice of parameters, Fig. 9 shows the population flux, $\eta$, as a function of environmental noise. We use the structures previously identified as optimal for short- and long-range chains of nine and ten sites, and examine the flux while varying the dephasing rate, $\Gamma$. Notably, a turnover in flux, a hallmark of ENAQT, appears only in structures optimized with long-range tunneling, while the flux in purely short-range tunneling systems shows no such peak. This is expected given the localization dynamics that develop in corrugated structures, and subsequently incoherent effects serve to enhance transport. The flux peak in the long-range case does not align exactly with the $\Gamma$ value used to obtain the optimized energy profile, since the optimization was performed at a fixed $\Gamma$ (indicated by the dashed line). However, for the nine- and ten-site chains, the long-range peaks lie close to this value of $\Gamma$, confirming that the system operates within the ENAQT regime.
In Appendix B, we repeat the optimization process of OQS Model I with shorter chains of five and six sites; see Fig. 19. Our findings for these shorter chains are consistent with the nine- and ten-site chains: for models with short-range tunneling, we obtain nearly flat energy profiles and a constant gradient population. In contrast, long-range tunneling benefits from non-monotonic, corrugated energy landscapes, with population profiles exhibiting a clear transition from a decaying trend to an almost constant distribution along the chain.
We summarize our key findings on optimizing carrier transport in chains subject to local dephasing with the following OQS I design rule: when local dephasing is strong and comparable in scale to the tunneling energies, the short-range tunneling model favors classical-like conduction with population decaying with length with a constant gradient. This is achieved through nearly flat energy landscapes with only a mild detuning from the entry and exit sites. In contrast, when long-range tunneling is allowed, optimal transport arises when the system successfully establishes quasi-ballistic conduction, as reflected by the steady-state population profiles. The observation of a relatively flat population profile at some sites away from the entrance ($n=4$ in Fig. 7(c)-(d) and Fig. 19(c)-(d)) under long-range tunneling is particularly striking. It demonstrates that efficient transport can emerge even in chains with highly non-uniform energy landscapes, provided that the energy structure is carefully balanced with the long-range tunneling amplitudes and environmental effects. This transport regime is enabled by a highly nontrivial energy landscape that promotes efficient end-to-end transfer.